[Ncep.nhc.nco_contacts] TSB Morning Rounds - Tue Aug 22, 2017

TSB Admin nhc.tsbadmin at noaa.gov
Tue Aug 22 15:02:15 UTC 2017


Here is today's summary of NHC computer operations:
--- Craig Mattocks

----------

TAFB:
1. GFE/NCP issues: TSB AWIPS II Sysadmin Chris Mello worked to get the full  
set of UK-MET model results into GFE/D2D; previously TAFB was not receiving  
all hours/timesteps.

2. N-AWIPS issues: NCO Linux Sysadmin Doug Gaer removed one of the two 1 GB  
graphics cards from the N-AWIPS workstation sheephead and replaced the  
remaining one with a 2 GB graphics card. TAFB needs to test loading large  
SPFs to see if this amount of video RAM is sufficient.

3. The 10z Experimental Pacific Offshores products OFFPZ7 and OFFPZ8 were  
not done/sent so there was a 12-hour gap between issuances. The 00z Pacific  
NWPS was run twice.
HSU:
1. The in-house FSU/Genesis tool is still not functioning correctly. The  
Aug 21/12Z UK was not available and both the Summary and Consensus were  
only available from Aug 21/00Z and not 12Z for both the NE Pacific and  
Atlantic basins. TSB/JHT Developer Jose Salazar is investigating.

2. NESDIS used the ABI MESO-1 GOES-16 product between 1645-2000 UTC  
yesterday to monitor the solar eclipse. ABI MESO-2 was still available for  
weather needs. NESDIS Environmental Satellite Processing Center (ESPC)  
approved NHC's request to run from 1900 UTC yesterday until 0300 UTC today.  
NHC's Chris Landsea called ESPC and was able to get coverage extended until  
21Z today for Harvey remnant monitoring.
Other items:
1. Comcast cable service was restored to WFO-MFL and NHC around 10:45 am  
Monday. If there are any CTV boxes in the building not working properly,  
please let Patty know so Comcast can service those.
NHC RFCs:


-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170822/44f9d5e9/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list